At the beginning of 2017, your company buys a $28,400 piece of equipment that it expects to use for 4 years. The equipment has an estimated residual value of 2,000. The company expects to produce a total of 200,000 units. Actual production is as follows: 45,000 units in 2017, 54,000 units in 2018, 51,000 units in 2019, and 50,000 units in 2020.


Required:

  1. Determine the depreciable cost.
  2. Calculate the depreciation expense per year under the straight-line method.
  3. Use the straight-line method to prepare a depreciation schedule.
  4. Calculate the depreciation rate per unit under the units-of-production method.
  5. Use the units-of-production method to prepare a depreciation schedule.
Answer #1

Answers

  • Requirement ‘a’

A

Cost

$            28,400.00

B

Residual Value

$              2,000.00

C=A - B

Depreciable Cost

$            26,400.00

  • Requirement ‘b’

A

Cost

$            28,400.00

B

Residual Value

$              2,000.00

C=A - B

Depreciable base

$            26,400.00

D

Life [in years]

4

E=C/D

Depreciation expense per year

$              6,600.00

  • Requirement ‘c’

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

2017

$              28,400.00

$              6,600.00

$         21,800.00

$                6,600.00

2018

$              21,800.00

$              6,600.00

$         15,200.00

$             13,200.00

2019

$              15,200.00

$              6,600.00

$            8,600.00

$             19,800.00

2020

$                8,600.00

$              6,600.00

$            2,000.00

$             26,400.00

  • Requirement ‘d’

A

Cost

$            28,400.00

B

Residual Value

$              2,000.00

C=A - B

Depreciable base

$            26,400.00

D

Usage

                   200,000

E = C/D

Depreciation rate per unit

$                       0.13

  • Requirement ‘e’

Year

Book Value

Usage

Depreciation expense

Ending Book Value

Accumulated Depreciation

2017

$              28,400.00

                     45,000

$            5,940.00

$             22,460.00

$           5,940.00

2018

$              22,460.00

                     54,000

$            7,128.00

$             15,332.00

$        13,068.00

2019

$              15,332.00

                     51,000

$            6,732.00

$                8,600.00

$        19,800.00

2020

$                8,600.00

                     50,000

$            6,600.00

$                2,000.00

$       26,400.00